# Sparse Attention Vectors: Generative Multimodal Model Features Are Discriminative Vision-Language Classifiers

### v2

@@ Abstract
Generative Large Multimodal Models (LMMs) like LLaVA and Qwen-VL excel at a wide variety of vision-language (VL) tasks such as image captioning or visual question answering. Despite strong performance, LMMs are not directly suited for foundational discriminative vision-language tasks (i.e., tasks requiring discrete label predictions) such as image classification and multiple-choice VQA. One key challenge in utilizing LMMs for discriminative tasks is the extraction of useful features from generative models. To overcome this issue, we propose an approach for finding features in the model’s latent space to more effectively leverage LMMs for discriminative tasks. Toward this end, we present Sparse Attention Vectors (SAVs)—a finetuning-free method that leverages sparse attention head activations (fewer than 1% of the heads) in LMMs as strong features for VL tasks. With only few-shot examples, SAVs demonstrate state-of-the-art performance compared to a variety of few-shot and finetuned baselines on a collection of discriminative tasks. Our experiments also imply that SAVs can scale in performance with additional examples and generalize to similar tasks, establishing SAVs as both effective and robust multimodal feature representations. Code: https://chancharikmitra.github.io/SAVs_website/.
footnotetext: *Denotes Equal Contribution.
@@ 1 Introduction
Large Multimodal Models (LMMs) such as GPT-4V [67], LLaVA [55, 54], and QwenVL [3] demonstrate state-of-the-art performance on open-ended vision-language (VL) tasks like image captioning [51, 99], visual question answering [2, 28, 39], and language grounding [34, 60]. However, despite their remarkable performance on generative tasks, these models struggle on discriminative tasks, where responses are restricted to a discrete set of labels [103, 7]. Indeed, LMMs with billions of parameters and trained on trillions more tokens of data underperform smaller discriminative VLMs [103, 7] and even classical machine learning methods [5] on image classification tasks. Nevertheless, there are many discriminative tasks that generative LMMs may be better suited for than CLIP-like models such as hallucination detection and VQA. Thus, it is enticing to have one type of model that can effectively accomplish both generative and discriminative vision-language tasks.
@@
Figure 1: Sparse Attention Vectors (SAVs) Overview. We develop a method for extracting features from a generative LMM without finetuning. We first extract a sparse set of attention vectors for each task given a set of few-shot examples, and then, we utilize these attention vectors directly as features for downstream discriminative vision-language tasks.
@@
Another drawback of generative LMMs is that it is still unclear how best to extract features directly as in discriminative VLMs like CLIP or SigLIP. Feature extraction is a well-explored field in both vision-only [87, 77, 12] and language-only discriminative models [11, 75], but such is not the case generative models. Most current methods for extracting features from generative models require carefully constructed prompts [37], specialized architectures [48], and finetuning [59]. However, generative models still offer the promise of more flexible, truly multimodal features as compared to modality-specific features extracted from CLIP-like models. As such, we are motivated to extract multimodal features from a generative LMM without finetuning to be used for any discriminative VL task.
A natural question arises is how best to empower generative LMMs with discriminative capabilities. The simplest strategy is to use prompt engineering [96, 62] and few-shot prompting [6, 104] that guide the model to output class labels [9]. However, recent work [103] shows that prompting an LMM for discriminative tasks does not close the gap with discriminative VLMs. This result suggests a more direct approach of finetuning the LMM on discriminative tasks. While the finetuning approach appears to work [7, 103], it veils the key problem of requiring training-scale data for every new discriminative task. As LMMs are a combination of a strong encoder VLM and a large language model trained on internet-scale data, it should suggest the existence of a more efficient way to enable performant, generalizable discriminative capabilities in LMMs without finetuning.
@@
One source of inspiration for our method is long standing work in the field of neuroscience that suggests certain areas of the brain are reserved for specific tasks [33, 14] (i.e. functional specificity). Motivated by this idea, we refer to recent interpretability research that has focused on identifying specific heads in transformer-based models that correspond to particular tasks [66]. The most prominent of these methods is a line of work that looks to enhance vision-language capabilities using task vectors [20, 86, 23, 26], which are compact implicit representations of tasks encoded in the activations of a transformer model. While promising, these methods ultimately use these representations to augment a model’s generative capabilities. On the other hand, we seek to use feature representations directly as classifiers. Nevertheless, this intuition from interpretability informs our work on Sparse Attention Vectors (SAVs), which are sparse features in an LMMs activation space that can be directly exploited for few-shot discriminative reasoning.
@@
Our method has three steps: First, we extract features (called attention vectors) from the output of each head of the LMM for some few-shot labeled examples ($\approx 20$ per label). Second, we average these attention vectors over the examples in each class and evaluate their accuracy as centroids in a class centroid classifier. We then select the top 20 heads by classfication accuracy as our SAVs. In this way, we identify a very sparse set of attention vectors (less than 1% of the total number of heads) that can be used for discriminative tasks. Finally, we perform inference on the given task by doing a majority vote across this sparse set of attention vectors for each new query. This approach requires only few-shot examples at test-time to extract effective multimodal embeddings. An overview is shown in Figure 1.
@@
We summarize our main contributions as follows: (i) We introduce a novel method that yields a sparse set of attention vectors (less than 1%) for each individual task can serve as highly effective features for discriminative tasks; (ii) We demonstrate that our method can help close the gap with discriminative VLMs on classification tasks using only few-shot examples at test time; (iii) Our method surpasses zero-shot, few-shot, and LoRA fine-tuned baselines across multiple tasks (+7% improvement on average over LoRA on challenging benchmarks like BLINK [15], VLGuard [106], and NaturalBench [40]); (iv) We establish several advantageous properties of our approach, including strong generalization capabilities and favorable scaling characteristics.
@@ 2 Related Works
Figure 2: Sparse Attention Vectors (SAVs) Detailed View. Our method is broken into the following three parts: (1) Given a set of few shot examples to be classified by a frozen LMM, we extract attention vectors across all heads given a set of few-shot examples for each class. (2) These attention vectors are averaged across the set of examples for each class. For each head, we use these mean attention vectors as features for a classifier, resulting in a classification accuracy for each attention vector. The set of sparse attention vectors are selected by top-k classification accuracy. (3) Finally, we use these sparse attention vectors to directly classify new inputs via majority vote.
@@
Controllable Generation for Classification. Controllable text generation in LMMs has become an important area of research, aiming to guide model outputs to adhere to specific attributes or constraints. In particular, an important application of controllable generation is utilizing a generative LMM for discriminative tasks. One controllable generation method often used is test-time hard prompting [96, 6], where prompt engineering or few-shot examples guide the model to output the desired class labels [9, 81, 94, 97, 103]. Another similar technique is directly using the probability of a specific class label being generated by the LMM [52, 53], which is commonly used for many image-text retrieval tasks. Furthermore, soft prompting methods that finetune specific learnable tokens [38, 46] can also be a viable option for discriminative classification with LMMs. Apart from these approaches, one can directly instruction finetune the model [95] on labeled classification data [103, 5]. Another option is preference modeling, which feature methods like DPO and RLHF to align the model to accurately output class labels [73, 13, 68]. In contrast, our method is a finetuning-free approach that directly chooses class labels without the need for preference data.
Most related are works which show that internal representations of transformer models called task vectors [85, 19, 23, 26] (or function vectors) can encapsulate tasks outlined by ICL examples. In these works, such vectors are patched directly back in the model for generation. Going beyond previous work, however, we utilize a very sparse set of attention vectors directly as features for a discriminative task.
@@
Vision-Language Features. The study of feature extraction in deep learning is concerned with finding useful representations that can be applied to a diverse array of downstream tasks. Early development of embedding techniques include autoencoder methods [4, 76, 35, 57, 58], Word2Vec [61] and GloVe [71] which transformed inputs into computable vector representations. These methods were quickly followed up by similar works in NLP [11, 75, 63, 16] and computer vision [87, 77, 12]. More recently, methods like CLIP or SigCLIP [72, 43, 44, 100, 101, 47] explore the correlation between multiples modalities (primarily images and text) through contrastive learning or a sigmoid loss on image-text pair data. The value of such representations is their flexibility in being applied to a variety of downstream tasks [10, 88, 36, 74, 21, 29] and domains [102, 93, 50].
Extracting features from generative models is a notably more challenging problem as it is not immediately obvious where in the model architecture to extract a distilled representation from (unlike discriminative models). Nevertheless, this direction is attractive due to the potential flexibility of the embeddings. Some methods finetune encoder VLM models on synthetically generated data from generative LLMs [45, 89]. Another more direct approach is to finetune an LLM or LMM directly on classification and similarity tasks [103, 27]. A more efficient line of work finetunes encoder-decoder and decoder-only representations directly to better align modalities or tasks [65, 64, 59, 32]. Another area of finetuning-free approaches proposes prompting the model with a customized distillation prompt (e.g. “[TEXT] The meaning of the previous sentence in one word is:”) and then extracting a representation from the weights or activations of the model [31, 30, 56, 37]. Other methodologies require the use of more complex methods like mixture-of-experts models [49], a finetuned expert model [92], or LLM-based embedding reranking [17].
@@
To summarize, the current SOTA still faces the following challenges when extracting features from generative models: (1) being limited to modality-specific rather than truly multimodal features, (2) requiring finetuning of the model or embedding, (3) limited flexibility due to specialized prompts, and (4) relying on expert or multiple models. Our approach remedies each of these problems. SAVs yield effective multimodal embeddings (as opposed to modality-specific embeddings) without the need for any gradient-based finetuning—whether the LMM or the embedding itself. Furthermore, SAVs can be flexibly applied to a variety of discriminative VL tasks without any additional models.
Safety VQA Image Cls.
Model MHalu VLGuard BLINK Natural Bench EuroSAT Pets
Text Image Group
CLIP - - - - - - 64.0 88.1
SigLip - - - - - - 63.9 98.3
LLaVA-OV-7B-ZS 34.7 31.4 45.0 52.0 53.3 27.0 66.5 88.1
+MTV 37.3 32.9 44.5 56.2 58.0 30.7 65.5 88.5
+Last L 65.3 77.9 41.9 43.8 45.3 20.3 56.8 88.6
+All L 56.6 92.4 41.7 57.4 59.3 31.7 73.1 75.4
+LoRA 44.9 43.8 47.0 58.6 60.9 32.4 85.0 96.8
+SAVs 80.8 94.3 51.8 60.3 62.3 35.1 86.7 97.0
Qwen2-VL-7B-ZS 24.0 26.9 43.3 53.8 56.6 28.5 54.7 92.6
+MTV 32.3 21.9 41.9 54.8 57.3 29.7 52.3 91.7
+Last L 74.9 83.9 38.8 52.7 55.1 26.9 52.9 91.3
+All L 58.8 91.1 40.1 42.9 51.6 24.4 52.2 92.6
+LoRA 84.8 87.7 46.3 55.3 57.4 28.8 72.9 98.4
+SAVs 85.1 96.0 47.2 57.6 60.9 32.3 79.9 98.1
Table 1: Results evaluation on Safety, Visual Question Answering (VQA), and Classification benchmarks. The best result for each generative model is shown in bold and the second best in underline. We gray out discriminative VLM model results, which cannot be evaluating directly on tasks with interleaved image-text queries.
@@ 3 Methods
In this section, we outline our approach for using sparse attention vectors from the activation space of a transformer-based large multimodal model (LMM) as features for any discriminative VL task. The method consists of three main steps: (i) extracting the attention vectors from all attention heads in the model, (ii) identifying a sparse set of vectors based on their ability to consistently return the correct label for some support set of examples, and (iii) using these sparse features to classify new queries. We begin with a formal description of the transformer decoder LLM and its attention mechanism, followed by the detailed methodology for sparse attention vector selection and classification. A detailed view of our method is shown in Figure 2.
@@ 3.2 Sparse Attention Vectors
Step 2: Identifying Relevant Vectors. The central question is how to identify which attention vectors are naturally suited for the discriminative task at hand. We evaluate each vector’s discriminative ability by computing its performance under a nearest class centroid classifier.
@@
Next, we measure the discriminative ability of each head by its performance as follows:
@@
where $s_{l,m}(\cdot,\cdot)$ is defined as in Step 2. Our final class prediction counts the majority vote across all heads in $\mathcal{H}_{\text{SAV}}$:
@@
This approach reveals a surprising capability of LMMs: with just a few carefully selected attention heads ($|\mathcal{H}_{\text{SAV}}|\ll LH$), we can transform a generative language model into a lightweight vision-language classifier. This finding suggests that classification-relevant features naturally emerge within specific attention heads during model pretraining.
Figure 3: Head Visualization. We show the top-20 attention head locations for a given task in the top row. On the bottom row, we visualize the attention vector of few-shot examples for the top head of the given class with t-SNE clustering [22].

### v3

# Enhancing Few-Shot Vision-Language Classification with Large Multimodal Model Features

@@ Abstract
Generative Large Multimodal Models (LMMs) like LLaVA and Qwen-VL excel at a wide variety of vision-language (VL) tasks. Despite strong performance, LMMs’ generative outputs are not specialized for vision-language classification tasks (i.e., tasks with vision-language inputs and discrete labels) such as image classification and multiple-choice VQA. One key challenge in utilizing LMMs for these tasks is the extraction of useful features from generative LMMs. To overcome this, we propose an approach that leverages multimodal feature extraction from the LMM’s latent space. Toward this end, we present Sparse Attention Vectors (SAVs)—a finetuning-free method that leverages sparse attention head activations (fewer than 5% of the heads) in LMMs as strong feature representations. With only few-shot examples, SAVs demonstrate state-of-the-art performance compared to a variety of few-shot and finetuned baselines on a collection of vision-language classification tasks. Our experiments also imply that SAVs can scale in performance with additional examples and generalize to similar tasks, establishing SAVs as both effective and robust multimodal feature representations.
@@ 1 Introduction
Generative Large Multimodal Models (LMMs) such as GPT-4V [72], LLaVA [59, 58], and QwenVL [3] demonstrate state-of-the-art performance on open-ended vision-language (VL) tasks like image captioning [53, 105], visual question answering [2, 31, 42], and language grounding [37, 64]. However, despite their remarkable performance on generative tasks, these models struggle on vision-language classification tasks, where responses are a discrete set of labels [109, 7]. Indeed, LMMs with billions of parameters and trained on trillions more tokens of data underperform smaller encoder VLMs like CLIP and SigLIP [109, 7] and even classical machine learning methods [5] on image classification tasks. One reason that CLIP-like models excel at classification is that is straightforward to extract visual (or textual) features from visual (or textual) encoders. However, such models are not equipped to process joint vision-language inputs, as LMMs can. Our goal is to extract CLIP-like features from generative LMMs, enabling downstream discrete labeling tasks such as classification and multiple-choice VQA.
@@
Figure 1: Sparse Attention Vectors (SAVs) Overview. We develop a method for extracting features from a generative LMM without finetuning. We first extract a sparse set of attention vectors for each task given a set of few-shot examples, and then, we utilize these attention vectors directly as features for downstream vision-language classification tasks.
@@
Feature extraction is a well-explored field in both vision-only [93, 82, 14] and language-only encoder models [12, 80], but such is not the case with generative models. Most current methods for extracting features from generative models require carefully constructed prompts [40, 102, 66], specialized architectures [50], few-shot prompting [6, 110, 10], and finetuning [63, 7, 109]. Recent work [109] shows that prompting approaches fail to close the gap with encoder VLMs, while finetuning requires training-scale data for every unseen task, which is inefficient. However, generative models still offer the promise of more flexible, truly multimodal features as compared to modality-specific features extracted from CLIP-like models. As such, we are motivated to extract multimodal features from a (frozen) generative LMM, for use in any downstream vision-language classification task.
@@
One source of inspiration for our method is long-standing work in the field of neuroscience that suggests certain areas of the brain are reserved for specific tasks [36, 16] (i.e. functional specificity). Motivated by this idea, we refer to recent interpretability research that has focused on identifying specific heads in transformer-based models that correspond to particular tasks [71]. The most prominent of these methods is a line of work that looks to enhance vision-language capabilities using task vectors [23, 92, 26, 29], which are compact implicit representations of tasks encoded in the activations of a transformer model. While promising, these methods ultimately use these representations to augment a model’s generative capabilities. On the other hand, we seek to use feature representations directly as classifiers. Nevertheless, this intuition from interpretability informs our work on Sparse Attention Vectors (SAVs), which are sparse features in an LMMs activation space that can be directly exploited for few-shot VL classification.
@@
Our method has three steps: First, we extract features (i.e., attention vectors) from the output of each head of the LMM for some few-shot labeled examples ($\approx 20$ per label). Second, we average these attention vectors over the examples in each class and evaluate their accuracy as centroids in a class centroid classifier. We then select the top 20 heads by classification accuracy as our SAVs. In this way, we identify a very sparse set of attention vectors (less than 5% of the total number of heads) that can be used for discriminative tasks. Finally, we perform inference on the given task by doing a majority vote across this sparse set of attention vectors for each new query. This approach requires only few-shot examples at test-time to extract effective multimodal embeddings. An overview is shown in Figure 1.
@@
We summarize our main contributions as follows: (i) We introduce a novel method that yields a sparse set of attention vectors to serve as highly effective features for individual VL classification tasks; (ii) Our method surpasses zero-shot, few-shot, and LoRA fine-tuned baselines across multiple tasks (+7% improvement on average over LoRA on challenging benchmarks like BLINK [17], VLGuard [112], and NaturalBench [43]); (iii) We establish several advantageous properties of our approach, including strong generalization capabilities and favorable scaling characteristics.
Figure 2: Sparse Attention Vectors (SAVs) Detailed View. Our method is broken into the following three parts: (1) Given a set of few shot examples to be classified by a frozen LMM, we extract attention vectors across all heads and layers. (2) These attention vectors are averaged across the set of examples for each class. For each head, we use these mean attention vectors as centroid classifiers, which are then used to select a sparse set of $k$ heads with the highest classification accuracy. (3) Finally, we use these sparse attention vectors to directly classify new inputs via majority vote.
@@ 2 Related Works
Controllable Generation for Classification. Controllable text generation in LMMs guides model outputs to meet specific constraints. For classification tasks with generative LMMs, several approaches exist: test-time hard prompting [102, 6] uses prompt engineering or few-shot examples to elicit class label outputs [10, 87, 100, 103, 109]; direct probability analysis of generated class labels [54, 57] for image-text retrieval; and soft prompting methods that finetune learnable tokens [41, 48]. Other techniques include instruction finetuning [101] on labeled data [109, 5] and preference modeling like DPO and RLHF [78, 15, 73]. Our method, however, is finetuning-free and directly selects class labels without preference data. Most related work shows that internal transformer representations called task vectors [91, 22, 26, 29] can encapsulate ICL example tasks. Beyond previous approaches, we use a sparse set of attention vectors directly as VL classification task task features.
@@
Vision-Language Features. Feature extraction seeks useful representations for diverse downstream tasks. Early embedding techniques including autoencoders [4, 81, 38, 61, 62], Word2Vec [65] and GloVe [76] transformed inputs into vector representations, followed by advances in NLP [12, 80, 67, 18] and vision [93, 82, 14]. Recent methods like CLIP and SigCLIP [77, 45, 46, 106, 107, 49] explore multimodal correlations through contrastive learning or sigmoid loss. These representations offer flexibility across tasks [11, 94, 39, 79, 24, 32] and domains [108, 99, 52].
Interestingly, extracting features from generative models poses unique challenges in identifying optimal extraction points. Some approaches finetune encoder VLMs on LLM-generated data [47, 96] or finetune LMMs directly on specific tasks [109, 30]. More efficient methods finetune encoder-decoder representations for better alignment [69, 68, 63, 35]. Finetuning-free approaches use distillation prompts to extract representations from model weights or activations [34, 33, 60, 40], while other methods employ mixture-of-experts [51], expert models [98], or embedding reranking [20].
@@
Thus, current SOTA faces the following challenges: (1) modality-specific rather than multimodal features, (2) requiring finetuning, (3) limited flexibility due to specialized prompts, and (4) dependence on multiple models. Our approach provides effective multimodal embeddings without any gradient-based finetuning and flexibly apply to various VL classification tasks without additional models.
@@ 3 Methods
In this section, we outline our approach for using sparse attention vectors from the activation space of a transformer-based large multimodal model (LMM) as features for any VL classification task. The method consists of three main steps: (i) extracting the attention vectors from all attention heads in the model, (ii) identifying a sparse set of vectors based on their ability to consistently return the correct label for some support set of examples, and (iii) using these sparse features to classify new queries. We begin with a formal description of the transformer decoder LLM and its attention mechanism, followed by the detailed methodology for sparse attention vector selection and classification. A detailed view of our method is shown in Figure 2.
@@ 3.2 Sparse Attention Vectors
Step 2: Identifying Relevant Vectors. The central question is how to identify which attention vectors are naturally suited for the classification task at hand. We evaluate each vector’s ability as a local classifier by computing its performance under a nearest class centroid classifier.
@@
Next, we measure the classification accuracy of each head (i.e. local classifier):
@@
where $s_{l,m}(\cdot,\cdot)$ is defined as in Step 2. Our final class prediction is that of a global classifier that counts the majority vote across all local classifiers (heads) in $\mathcal{H}_{\text{SAV}}$:
@@
This approach reveals a surprising capability of LMMs: with just a few carefully selected attention heads ($|\mathcal{H}_{\text{SAV}}|\ll LH$), we can transform a generative language model into a lightweight vision-language classifier. This finding suggests that classification-relevant features naturally emerge within specific attention heads during model pretraining. In theory, one could replace our local and global classifiers with more complex varients (e.g. linear models, MLPs), something we explore in 3(a).
Image-Text Tasks Image-Only Tasks
Safety VQA I-T Retrieval Classification
Model MHalu VLGuard BLINK VizWiz NB (T) NB (I) NB (G) NB-r (T) NB-r (I) NB-r (G) SC EuroSAT Pets Imagenet-1k Flower CUB
CLIP - - - - - - - 41.8 45.0 23.2 35.3 64.0 88.1 96.1 92.8 97.8
SigLip - - - - - - - 54.5 54.9 31.2 42.7 63.9 98.3 97.6 95.8 98.4
GPT-4o 39.4 45.7 59.0 58.1 64.6 66.4 39.6 65.0 67.0 40.5 48.3 70.5 98.3 98.3 99.0 99.3
LLaVA-1.5 44.3 31.0 38.0 50.0 37.7 43.8 12.7 36.7 42.7 12.2 24.9 26.3 43.0 20.6 51.8 49.0
Instruct-BLIP 44.0 18.5 38.7 34.5 20.2 24.2 4.0 19.5 21.3 1.1 4.7 20.4 56.0 11.0 18.2 33.8
LLaVA-OV-7B 34.7 31.4 45.0 60.4 52.0 53.3 27.0 56.2 58.0 32.1 15.3 66.5 88.1 92.5 83.2 85.3
+4-shot-ICL 25.0 35.0 38.9 47.8 47.6 50.4 22.2 48.2 49.6 31.5 16.1 47.1 63.9 49.0 63.8 60.6
+MTV 37.3 32.9 44.5 61.1 56.2 58.0 30.7 58.1 59.5 33.4 28.7 65.5 88.5 95.9 83.2 85.6
+LoRA 78.3 90.0 47.0 63.1 58.6 60.9 32.4 59.4 60.3 35.4 30.4 85.0 96.8 93.2 91.2 91.8
+SAVs 80.8 94.3 51.8 66.1 60.3 62.3 35.1 72.7 73.0 53.1 37.6 86.7 97.0 97.5 99.6 97.5
+46.1 +62.9 +6.8 +5.7 +8.3 +9.0 +8.1 +16.5 +15.0 +20.5 +22.3 +20.2 +8.9 +5.0 +16.4 +12.2
Qwen2-VL-7B 24.0 26.9 43.3 68.3 53.8 56.6 28.5 60.2 61.9 35.6 24.9 54.7 92.6 84.4 93.7 93.2
+4-shot-ICL 40.4 52.9 37.6 67.1 38.2 41.3 15.2 42.4 45.6 22.7 25.2 29.4 43.5 43.8 59.4 48.4
+MTV 32.3 21.9 41.9 68.5 54.8 57.3 29.7 63.5 64.0 37.0 40.7 52.3 91.7 91.2 94.3 94.1
+LoRA 84.8 87.7 46.3 70.8 55.3 57.4 28.8 65.2 66.1 40.4 38.4 72.9 98.4 96.4 97.1 95.0
+SAVs 85.1 96.0 47.2 68.3 57.6 60.9 32.3 70.0 71.0 42.5 47.5 79.9 98.1 97.6 99.8 98.7
+61.1 +69.1 +3.9 +0 +3.8 +4.3 +3.8 +9.8 +9.1 +13.8 +22.6 +25.2 +5.5 +13.2 +6.1 +5.5
Table 1: Results evaluation on Safety, Visual Question Answering (VQA), Image-Text Retrieval (I-T Retrieval), and Image Classification benchmarks. The best result for each generative model is shown in bold and the second best in underline. We gray out additional baselines. We note that CLIP and SigLIP cannot be evaluating directly on tasks with interleaved image-text queries. $\dagger$ GPT-4o is a close-source model and as such, is shown just as an upperbound (since SAVs are not directly applicable). Key: NB - NaturalBench, SC - SUGARCREPE.